# Learning production functions for supply chains with graph neural networks

### v1

@@ 2.3 Model training and evaluation
Our model does not have access to any production functions during training, but the goal is for the model to learn production functions via the inventory module and its specialized loss. For each product $p$, we have the set of its attention weights $\{\alpha_{pp_{1}},\alpha_{pp_{2}},\cdots,\alpha_{pp_{m}}\}$, and we can compute the ranking over all products from highest to lowest weight. We compare this ranking to the product’s ground-truth parts, i.e., the parents of $p$ in $G_{\textrm{prod}}$, and use average precision (17) to quantify performance. Then, we compute the mean average precision (MAP) over all products for which we have ground-truth parts. We have ground-truth parts for all products in our synthetic data (Section 3.2) and for scanning electron microscopes in real-world data (Section 3.1).
@@
We perform negative sampling such that each positive transaction $e(s,b,p,t)$ is paired with a set of negative transactions that did not actually occur at time $t$. Following Temporal Graph Benchmark (Huang et al.,, 2023), we sample two types of hard negatives: first, randomly perturbing one of the three nodes; second, sampling a historical negative, meaning a transaction that appeared in training but not at time $t$. Compared to a uniform random negative, these negatives are far harder to distinguish from the positive transaction. For the perturbations, they match the positive transaction on two of the three nodes, requiring the model to make a distinction based on the third node. The historical negatives are even harder: this transaction has been observed before, but it does not occur at this timestep, requiring the model to learn time-varying properties of the graph. For each positive transaction, we sample 9 perturbation negatives and 9 historical negatives. To evaluate performance, we use mean reciprocal rank (MRR), which evaluates the rank of the positive transaction among the negatives (19). However, MRR is non-differentiable, so during training, we use softmax cross-entropy as a strong proxy loss for MRR (Bruch et al.,, 2019).
@@ 3 Supply chains data
To evaluate our models, we collected two real-world datasets of supply chain transactions, and built a simulator, SupplySim, which can generate realistic supply chain data under varied settings. While we cannot release the real-world datasets, due to their proprietary nature, we carefully document our data sources and collection processes. Furthermore, we release SupplySim (which matches the real data on key characteristics) and the three synthetic datasets that we evaluate our models on in this work. We describe both data types here, with details in Appendix B and data statistics in Table 3.
@@ 3.1 Real-world supply chains data
We acquired transactions data from TradeSparq, a third-party data provider which aggregates data from authorized government sources (e.g., customs departments) across 60+ countries. Their data sources include bills of lading, receipts of reported transactions, and custom declarations. Each product being transacted is described with a Harmonized System (HS) code, where HS6 is an internationally recognized six digit number that describes products in terms of their chapters (first two digits), headings (next two digits), and subheadings (last two digits). For example, the HS code 840731 refers to the chapter 84 (chapter: Machinery), 07 (heading: Spark-ignition Reciprocating Piston Engines), and 31 (sub-heading: engines type not over 50CC cylinder capacity). Along with HS codes, the transactions data also includes the company name and ID of the supplier, the company name and ID of the buyer, timestamp of the transaction, cost in USD, and so on. Using the TradeSparq data (which we could only access via API calls, with costs per call), we constructed two supply chains datasets. For both datasets, we began by selecting specific HS codes that we were interested in, identifying their prominent suppliers, then using breadth-first-search to collect their neighbors upstream and downstream in the supply chain (see flowchart in Figure 3). This collection process allowed us to capture all interactions between firms within multiple tiers of the supply chain.
Automotive dataset (Tesla). This dataset focuses on electric vehicles (EV) and EV parts supplied by Tesla. We identified Tesla EV makers, their direct suppliers and buyers, and their suppliers’ suppliers, and included all transactions between these firms from January 1, 2019, to December 31, 2022.
Industrial equipment dataset (IED). This dataset focuses on microscopes along with other specialized analytical and inspection equipment and their manufacturers. We included makers of these products and their direct suppliers and buyers, and included all transactions between these firms in 2023. Compared to the Tesla dataset, this dataset is broader (i.e., more firms per layer) but shorter in time and depth (i.e., fewer layers). For this dataset, we have also estimated the ground-truth parts of microscopes in terms of their HS codes, which we use to test our inventory module’s ability to infer production functions from transactions.
@@ 3.2 Supply chains simulator, SupplySim
The TradeSparq data offers a rare opportunity to test models on real transaction-level data, but it also has shortcomings: the transactions are incomplete, most production functions are missing, and we cannot release the data. Thus, we design a simulator, SupplySim, that addresses these shortcomings and enables us to test the model under controlled settings (e.g., complete vs. incomplete transactions, stable supply vs. shocks). There are three stages to the simulator: (1) constructing the production graph, $G_{\textrm{prod}}$, (2) assigning products to suppliers and assigning suppliers to buyers, (3) generating time-varying transactions between firms. We ensure that our generated graphs match real-world supply chain networks on known characteristics, such as power law degree distributions.
Constructing the production graph, $G_{\textrm{prod}}$.
Given the number of products, $m$, we partition the products into tiers and sample a 2-dimensional position for each product from Uniform($0,1$). Then, for each tier, we assign the products in the tier to parts from the previous tier, with probability proportional to the inverse distance between their positions. For each part-product pair, we also sample $u_{io}$, the number of units of the part $p_{i}$ needed to make one unit of the product $p_{o}$. The tier structure imitates tiers in real-world supply chains; for example, products at the first tier are raw materials, e.g., minerals, and products at the final tier are consumer products. The products’ positions capture the product type, e.g., its industry, and assigning parts based on positions reflects how parts should be similar to their products and naturally results in commonly co-occurring parts.
Constructing supplier-buyer relations.
Given the number of firms, $n$, we also sample a 2-dimensional position from Uniform($0,1$) for each firm, and we restrict the firm to a subset of two consecutive tiers, meaning it can only produce products in those tiers. Then, for each product, we select its suppliers from the set of firms that are allowed to produce that product, again with probability proportional to inverse distance. Now, each firm has a set of products that it is supplying, which means, based on $G_{\textrm{prod}}$, we know which input parts it needs to buy. For each pair $(b,p)$, where firm $b$ needs to buy product $p$, we assign it to a supplier of $p$ with probability proportional to the number of buyers that the supplier already has. This assignment mechanism, known as preferential attachment (Newman,, 2001), yields power law degree distributions, which are known to appear in real-world supply chain networks (Fujiwara and Aoyama,, 2010; Zhao et al.,, 2019).
Generating transactions.
We generate transactions based on the ARIO model, an agent-based model widely used in economics to simulate propagation over supply chains (Hallegatte,, 2008; Inoue and Todo,, 2019; Guan et al.,, 2020). At each timestep of the simulation, each firm completes as many of its incomplete orders as it can until it runs out of inventory. Each time it completes an order for $k$ amount of output product $p_{o}$, it subtracts $k\cdot u_{io}$ of input product $p_{i}$ from its inventory (i.e., we assume the production function $\mathcal{F}_{p_{o}}$ is linear). At the end of the timestep, the firm places orders to its suppliers, based on what it needs to complete its remaining orders. Finally, the reported transactions are the completed orders in each timestep. The simulator also keeps track of the exogenous supply for products in the first tier, which do not require parts, and of exogenous demand for products in the final tier, which are only bought by consumers and not by other firms. By manipulating the exogenous supply, we can model shocks in the supply chain.
4 Experiments
We run experiments on the two real-world supply chain datasets and three synthetic datasets from SupplySim: a standard setting with high exogenous supply (“SS-std”), a setting with shocks to exogenous supply (“SS-shocks”), and a setting with missing transactions (“SS-missing”), where we sampled 20% of firms uniformly at random and dropped all transactions where they were buyer or supplier, reflecting firms that are missing from real-world supply chains data. In all experiments, we order the transactions chronologically and split them into train (the first 70%), validation (the following 15%), and test (the last 15%). Here we describe our results on these datasets, with additional details about the experimental set-up and supplemental results in Appendix C.
4.1 Learning production functions
@@
1) Temporal correlations: we expect that inputs and outputs are temporally correlated, so this method computes the maximum correlation, with possible lags, between the buying timeseries of $p_{i}$ and supplying timeseries of $p_{o}$, averaged over all firms that buy $p_{i}$ and supply $p_{o}$.
2) Pointwise Mutual Information (PMI): we expect that input-output pairs appear with greater frequency in the firm-product graph, [^3] The firm-product graph is a static bipartite graph of firm and product nodes, where an edge between a firm and product indicates that the firm buys or supplies the product. so this method computes the probability that a firm buys $p_{i}$ and supplies $p_{o}$, divided by the product of their individual probabilities (26).
3) node2vec: we expect that inputs are close to outputs, so this method computes the cosine similarity of $p_{i}$ and $p_{o}$’s node2vec embeddings, learned over the firm-product graph.
Figure 2 visualizes our results, showing that the inventory module effectively learns the true production functions. The block structure visible in all three subfigures reflects how our products were partitioned into tiers (products 0-4 in tier 0, 5-14 in tier 1, 15-24 in tier 2, and so on) and products in one tier served as viable parts for the next tier. The inventory module recovers the tier structure much more precisely, while temporal correlations show positive weights in larger blocks. The inventory module also learns attention weights of similar magnitude as the true production functions, while the temporal correlations are constrained between 0 and 1, and the other baselines are similarly not comparable to the true production functions on magnitude, only ranking. While only ranking matters for MAP, magnitude is essential if using the inventory module to inform edge prediction with penalties (8)-(9). Finally, in Figure 6, we show that our inventory module’s loss function (6) is well-correlated with MAP, which is why it can effectively learn production functions without observing any of them.
Figure 2: True production functions (left), predictions from inventory module (middle), predictions from temporal correlations (right), trained on SS-std.
4.2 Predicting future edges
1) Edgebank: the “binary” version predicts 1 if $(s,b,p,*)$ appeared before in the train set; 0 otherwise. The “count” version predicts the number of times that $(s,b,p,*)$ appeared in the train set.
2) Static: learns a static vector to represent each node (no GNN).
3) Graph transformer: learns a static embedding to represent each node, using the UniMP model (Shi et al.,, 2021) which utilizes a graph transformer architecture and multi-head attention.
While Edgebank simply memorizes the train set, it serves as a strong baseline, as the Temporal Graph Benchmark (Huang et al.,, 2023) found that even the binary version outperformed some GNNs on dynamic link prediction. The latter two baselines test two types of static node embeddings, which allow us to isolate the benefit of dynamic node embeddings. For each of our models (SC-TGN and SC-GraphMixer), we try them alone and with the inventory module (+inv). For edge weight, we test the same set of methods, except for Edgebank, where we now predict the average amount for $(s,b,p,*)$ in the train set, or 0 (which is the mean amount, after standard scaling) if the triplet never appeared in train.
@@ 4.2 Predicting future edges
Table 2: Results for edge prediction. Top 8 rows are edge existence, evaluated with mean reciprocal rank (MRR $\uparrow$). Bottom 7 rows are edge weight (i.e., transaction amount), evaluated with root mean squared error (RMSE $\downarrow$). We report mean and standard deviation (in parentheses) over 10 seeds.
@@
We summarize the edge prediction results in Table 2. First, we find that our models, SC-TGN and SC-GraphMixer, outperform the baselines on both tasks and all five datasets, by 11-62% on edge existence and 1-13% on edge weight. Second, we find that SC-TGN consistently outperforms SC-GraphMixer; we hypothesize this is because of SC-TGN’s sophisticated updating of node memories, while SC-GraphMixer only captures changes over time through its link encoder, which encodes the features and timestamps of the node’s recent edges, and its node encoder relies on the node’s static features (Appendix A.2). Supporting this hypothesis is the fact that the static methods perform poorly, demonstrating the need for TPGs to be represented with dynamic node embeddings. Since we find that SC-TGN outperforms SC-GraphMixer, we also compared SC-TGN to ablated versions: the original TGN, to test the value of our extensions (which we document in Appendix A.1), and to SC-TGN where the node memory is directly used as the embedding, instead of applying a GNN (UniMP in our case) to the memories. We find that the full SC-TGN outperforms both ablations substantially, by 34-45% and 53-100%, respectively (Table 5).
@@
We also find in the synthetic data experiments, the setting with shocks is the hardest, resulting in lower MRRs and higher RMSEs across models. This demonstrates how shocks complicate prediction on TPGs, since shocks both delay firms’ abilities to complete their orders and limit them to producing smaller amounts, thus affecting both edge existence and edge weight. However, adding the inventory module significantly improves MRRs for both SC-TGN and SC-GraphMixer, on standard synthetic data and with shocks. In general, even though adding the inventory module introduces an additional loss so that we can learn production functions (16), we find that it does not hurt edge prediction performance, and in several cases, improves performance.
5 Related work
Many real-world systems can be represented as temporal graphs, such as transportation systems (Jiang and Luo,, 2022), human mobility patterns (Chang et al.,, 2023), and biological networks (Prill et al.,, 2005). While most GNNs are designed for static graphs, in recent years, there has been growing interest in developing GNNs for temporal graphs, as noted by the Temporal Graph Benchmark (Huang et al.,, 2023) and surveys on temporal GNNs (Skarding et al.,, 2021; Longa et al.,, 2023). However, to the best of our knowledge, GNNs have not yet been designed to handle temporal production graphs (TPGs), as described in this work. Furthermore, only a handful of works have explored GNNs for supply chains: Aziz et al., (2021) and Kosasih and Brintrup, (2022) use GNNs to predict hidden links between firms in the static supply chain network, and Wasi et al., (2024) provide a benchmark dataset for GNNs, consisting of a graph where nodes are products and edges represent relationships between products (e.g., same product group, same storage location). In contrast, we construct a dynamic supply chain network consisting of firm and product nodes, and our goals are to learn production functions and to forecast future transactions on this network.
Our work also has connections to temporal causal discovery (Nauta et al.,, 2019; Löwe et al.,, 2022; Assaad et al.,, 2022), where the goal is to learn a causal graph between variables based on their timeseries. Here, the variables are the products, but our setting is different since the timeseries appear as in-edges and out-edges embedded within the supply chain network, with each product having multiple timeseries (for different firms that buy and supply it). Our work is also related to inferring networks from node marginals (Kumar et al.,, 2015; Maystre and Grossglauser,, 2017; Chang et al.,, 2024) and combining deep learning with domain-specific principles, a theme which appears in physics simulation (Wang et al.,, 2021) and epidemiological forecasting (Liu et al.,, 2024).
Within the supply chains domain, our work builds on prior literature that represents supply chains as networks (Fujiwara and Aoyama,, 2010) and models the propagation of shocks over these networks (Acemoglu et al.,, 2012; Zhao et al.,, 2019; Li et al.,, 2020; Carvalho et al.,, 2021). Our work is unique in two key ways: first, much of the prior work relies on theoretical models and synthetic networks, and, even among empirical studies, the data used is typically industry-level or, at best, firm-level with static relations between firms (Carvalho and Tahbaz-Salehi,, 2019). In contrast, we have access to transaction-level data, which reveals essential time-varying information. Second, integration of ML into supply chains has been limited, and only recently have researchers begun exploring the use of ML in supply chains, with calls for further exploration (Baryannis et al.,, 2019; Brintrup et al.,, 2020; Younis et al.,, 2022). Most prior models for modeling supply chain propagation are mechanistic (Hallegatte,, 2008; Guan et al.,, 2020; Inoue and Todo,, 2020; Li et al.,, 2021) and limited in their ability to even fit aggregate counts, such as country-level GDP (Inoue and Todo,, 2019). In contrast, by combining mechanistic models with GNNs, we can forecast individual transactions with remarkable accuracy, while maintaining the interpretability of the inventory module. Our model can also be used for critical supply chain applications, such as demand forecasting (Seyedan and Mafakheri,, 2020), early detection of risks (Sheffi,, 2015), and inventory optimization (Vandeput,, 2020).
6 Conclusion

### v3

@@ 3.3 Model training and evaluation
Our model does not have access to any production functions during training, but the goal is for the model to learn production functions via the inventory module and its specialized loss. For each product $p$, we have the set of its attention weights $\{\alpha_{pp_{1}},\alpha_{pp_{2}},\cdots,\alpha_{pp_{m}}\}$, and we compute the ranking over all products from highest to lowest weight. We compare this ranking to the product’s ground-truth parts and use average precision (17) to quantify performance. Then, we compute the mean average precision (MAP) over all products for which we have ground-truth parts.
@@
We perform negative sampling such that each positive transaction $e(s,b,p,t)$ is paired with a set of negative transactions that did not actually occur at time $t$. Following Temporal Graph Benchmark (Huang et al. 2023), we sample two types of hard negatives: first, randomly perturbing one of the three nodes; second, sampling a historical negative, meaning a transaction that appeared in training but not at time $t$. For each positive transaction, we sample 9 perturbation negatives and 9 historical negatives. To evaluate performance, we use mean reciprocal rank (MRR), which evaluates the rank of the positive transaction among the negatives (21). However, MRR is non-differentiable, so during training, we use softmax cross-entropy as a strong proxy loss for MRR (Bruch et al. 2019).
@@ 4.1 Real-world supply chains data
We acquired transactions data from TradeSparq, a third-party data provider which aggregates data from authorized government sources across 60+ countries. Their data sources include bills of lading, receipts of reported transactions, and customs declarations. Each product is described with a Harmonized System (HS) code, an internationally recognized system for classifying products. Along with HS codes, the transactions data also includes the supplier firm, buyer firm, timestamp of the transaction, cost in USD, and more. Using the TradeSparq API, we constructed two datasets:
  1. 1.
    Our Tesla dataset focuses on electric vehicles (EV) and EV parts supplied by Tesla. We identified Tesla EV makers, their direct suppliers and buyers, and their suppliers’ suppliers, and included all transactions between these firms from January 1, 2019, to December 31, 2022.
  2. 2.
    Our industrial equipment dataset (IED) focuses on microscopes along with other specialized analytical and inspection equipment and their manufacturers. We included makers of these products and their direct suppliers and buyers, and included all transactions between these firms in 2023. For this dataset, we have also estimated the ground-truth parts of microscopes in terms of their HS codes, which we use to test our inventory module’s ability to infer production functions from transactions.
@@ 4.2 Supply chains simulator, SupplySim
Figure 2: SupplySim generates data matching real data on key characteristics: (a) power law degree distribution, (b) community structure, (c) low clustering, (d) time-varying transactions, with possible shocks or missing data.
@@
The TradeSparq data offers a rare opportunity to test models on real transactions data, but it also has shortcomings: we cannot release the data, it does not include most production functions, and, even though it provides more detailed information than most supply chains datasets, its transactions are still incomplete (e.g., missing domestic transactions). Thus, we design a simulator, SupplySim, that addresses these shortcomings and enables us to test the model under controlled settings (e.g., how much data is missing).
To ensure realism, our simulator incorporates many real-world aspects of supply chains: for example, firms specialize in certain tiers of products and time-varying transactions are generated based on the commonly used ARIO model from economics (Hallegatte 2008), which describes how firms complete orders from buyers and place orders to suppliers. We also show that our synthetic data matches real supply chain networks on key characteristics (Figure 2). For example, there is community structure, but fewer triangles (thus, lower average clustering coefficient), since, unlike social networks, a firm’s supplier’s supplier is unlikely to also be this firm’s supplier (Fujiwara and Aoyama 2010; Zhao, Zuo, and Blackhurst 2019). Our synthetic data also exhibits power law degree distributions, known to appear in real networks, and time-varying transactions with possible shocks or missing data. Below, we briefly describe the steps of our simulator, with details in Appendix B.2 and B.3.
Constructing the production graph, $G_{\textrm{prod}}$.
First, we partition the products into tiers (e.g., products 0-4 in tier 0, 5-14 in tier 1, 15-24 in tier 2, etc.) and sample a 2-dimensional position for each product from Uniform($0,1$). For each tier, we assign the products in the tier to parts from the previous tier, with probability proportional to the inverse distance between their positions. For each part-product pair, we sample $u_{io}$, the number of units of part $p_{i}$ needed to make one unit of product $p_{o}$. The tier structure imitates tiers in real supply chains, from raw materials in the first tier to consumer products in the final tier. The products’ positions capture the product type, e.g., its industry, and assigning parts based on positions reflects how parts should be similar to their products and naturally results in commonly co-occurring parts.
Constructing supplier-buyer graph.
For each firm, we also sample a 2-dimensional position from Uniform($0,1$), and we restrict it to two consecutive tiers, meaning it can only produce products in those tiers. Then, for each product, we select its suppliers from the set of firms that are allowed to produce that product, again with probability proportional to inverse distance. Now, each firm has a set of products that it is supplying, which means, based on $G_{\textrm{prod}}$, we know which input parts it needs to buy. For each pair $(b,p)$, where firm $b$ needs to buy product $p$, we assign it to a supplier of $p$ with probability proportional to the number of buyers that the supplier already has. This assignment mechanism, known as preferential attachment (Newman 2001), yields power law degree distributions (Figure 2a) known to appear in real supply chain networks (Fujiwara and Aoyama 2010; Zhao, Zuo, and Blackhurst 2019).
Generating transactions.
We generate transactions based on the ARIO model, an agent-based model widely used in economics to simulate propagation over supply chains (Hallegatte 2008; Inoue and Todo 2019; Guan et al. 2020). At each timestep of the simulation, each firm completes as many of its incomplete orders as it can until it runs out of inventory. At the end of the timestep, the firm places orders to its suppliers, based on what it needs to complete its remaining orders. Finally, the reported transactions are the completed orders in each timestep. The simulator also keeps track of the exogenous supply for products in the first tier, which do not require parts, and of exogenous demand for products in the final tier, which are only bought by consumers and not by other firms. By manipulating the exogenous supply, we can model shocks in the supply chain.
5 Experiments
We run experiments on the two real supply chain datasets and three synthetic datasets from SupplySim: a standard setting with high supply (“SS-std”), a setting with shocks to supply (“SS-shocks”), and a setting with missing transactions (“SS-missing”), where we sampled 20% of firms uniformly at random and dropped all of their transactions (Figure 2d). In all experiments, we order the transactions chronologically and split them into train (the first 70%), validation (the following 15%), and test (the last 15%). Here we describe our results on these datasets, with additional experimental details and results in Appendix C.
5.1 Learning production functions
@@
Table 2: Results for predicting existence of future edges, evaluated with mean reciprocal rank (MRR $\uparrow$). See Table 5 for edge weight results. We report mean and standard deviation (in parentheses) over 10 seeds.
  1. 1.
    Temporal correlations: we expect that inputs and outputs are temporally correlated, so this method computes the maximum correlation, with possible lags, between the buying timeseries of $p_{i}$ and supplying timeseries of $p_{o}$, averaged over all firms that buy $p_{i}$ and supply $p_{o}$.
  2. 2.
    Pointwise Mutual Information (PMI): we expect that input-output pairs appear with greater frequency in the firm-product graph, [^3] The firm-product graph is a static bipartite graph of firm and product nodes, where an edge between a firm and product indicates that the firm buys or supplies the product. so this method computes the probability that a firm buys $p_{i}$ and supplies $p_{o}$, divided by the product of their individual probabilities (29).
  3. 3.
    node2vec: we expect that inputs are close to outputs, so this method computes the cosine similarity of $p_{i}$ and $p_{o}$’s node2vec embeddings from the firm-product graph.
Figure 3: True production functions (left) and predictions from inventory module (right), trained on SS-std.
Figure 3 visualizes our results, showing that the inventory module effectively learns the true production functions. The inventory module also learns attention weights of similar magnitude as the true production functions, while the other baselines are not comparable on magnitude, only ranking. While only ranking matters for MAP, magnitude is essential if using the inventory module to inform edge prediction with penalties (14)-(15). Finally, in Figure 5, we show that our inventory module’s loss function (6) is well-correlated with MAP, which is why it can effectively learn production functions without observing any of them.
5.2 Predicting future edges
  1. 1.
    Edgebank: “binary” predicts 1 if $(s,b,p,*)$ appeared before in the train set; 0 otherwise. “count” predicts the number of times that $(s,b,p,*)$ appeared in the train set.
  2. 2.
    Static: learns a static vector to represent each node.
  3. 3.
    Graph transformer: learns a static embedding to represent each node, using the graph transformer model called UniMP from Shi et al. (2021).
While Edgebank simply memorizes the train set, it serves as a strong baseline, as the Temporal Graph Benchmark (Huang et al. 2023) found that even the binary version outperformed some GNNs on dynamic link prediction. The latter two baselines test two types of static node embeddings, which allow us to isolate the benefit of temporal GNNs. For each of our models (SC-TGN and SC-GraphMixer), we try them alone and with the inventory module (+inv).
@@ 5.2 Predicting future edges
In Table 2, we summarize results for predicting edge existence; results for predicting edge weight tell a very similar story (Table 5). First, we find that our models, SC-TGN and SC-GraphMixer, outperform the baselines on both tasks and all five datasets, by 11-62% on edge existence and 1-13% on edge weight. Second, we find that SC-TGN consistently outperforms SC-GraphMixer; we hypothesize this is because of SC-TGN’s sophisticated updating of node memories, while SC-GraphMixer only captures changes over time through its link encoder, which encodes the features and timestamps of the node’s recent edges, and its node encoder relies on the node’s static features (Appendix A.2). Supporting this hypothesis is the fact that the static methods perform poorly, demonstrating the need for temporal GNNs for supply chains. Since we find that SC-TGN outperforms SC-GraphMixer, we also compared SC-TGN to ablated versions: the original TGN, to test the value of our extensions (which we document in Appendix A.1), and to SC-TGN where the node memory is directly used as the embedding, instead of applying a GNN to the memories. We find that the full SC-TGN outperforms both ablations substantially, by 34-45% and 53-100%, respectively (Table 7).
@@
We also find in the synthetic data experiments, the setting with shocks is the hardest for all models. This demonstrates how shocks complicate prediction, since shocks both delay firms’ abilities to complete their orders and limit them to producing smaller amounts, thus affecting both edge existence and edge weight. However, adding the inventory module significantly improves both SC-TGN and SC-GraphMixer’s MRRs under shocks. In general, even though adding the inventory module introduces an additional loss so that we can learn production functions (16), we find that it does not hurt edge prediction performance, and in several cases, improves performance.
5.3 Case study: generating future transactions under shocks
In this section, we demonstrate a possible use case of our model: generating future transactions under supply chain shocks and subsequent disruptions. For this case study, we use the synthetic data from SS-shocks as our ground truth, since it specifically simulates disruptions propagating over the supply chain network, due to shocks in the supply of the Tier 0 products. We test whether our SC-TGN model (with inventory), which was trained on the first 70% of timesteps in SS-shocks, can generate transactions for the first $k$ timesteps after the training data ends. The last timestep in the training data from SS-shocks is $t=138$; as shown in Figure 2d, this timestep falls in the middle of a disruption, as the total number of transactions is falling around this timestep.
To enable this case study, we need to extend our model so that it can generate future transactions for multiple timesteps. We do this by utilizing its existing capabilities in predicting edge existence and edge weight:
  1. 1.
    Predict whether a transaction exists at timestep $t$, by using the decoder (7) to provide a score $\hat{y}$ for each possible transaction then keeping all transactions above score threshold, i.e., $\hat{y}\geq\tau$,
  2. 2.
    Conditioned on a transaction existing, predict its amount,
  3. 3.
    Use the predicted transactions and amounts (as opposed to the ground-truth transactions) to update the model and inventories before timestep $t+1$.
We use the training data to set the threshold $\tau$, by gathering the trained model’s edge existence scores over all positive and negative transactions in the training data (using the negative samples described in Section 3.3) and setting $\tau$ such that the number of scores greater than or equal to $\tau$ matches the number of positive transactions in the training data.
$t$ AUROC (all) AUROC (train)
139 0.996 0.824
140 0.986 0.733
141 0.966 0.721
142 0.938 0.637
143 0.943 0.637
144 0.936 0.593
145 0.936 0.634
Table 3: Generating transactions after the end of the SS-shocks training data ($t=138$) and evaluating the transaction existence scores using AUROC by comparing positive transactions to all negatives (all) and train negatives (train).
Using this approach, we generate transactions for $t=\{139,140,\cdots,145\}$, i.e., seven timesteps after the end of the training data. At each timestep $t$, we evaluate how well the predicted edge existence scores can discriminate between the ground-truth positive transactions vs. negative transactions that did not occur at timestep $t$. We compare to two types of negative transactions: all negatives, which keep all triplets that do not appear at time $t$, and train negatives, which keep all triplets appeared as transactions in the training data but did not appear at time $t$. While we already evaluated the model’s ability to predict edge existence in Table 2, this is a harder setting, since we are trying to forecast transactions for several timesteps without observing any ground-truth data, which means that the model is updated with its own predicted transactions and has the risk of propagating errors over multiple timesteps.
In Table 3, we report the area under the ROC curves (AUROC) of the predicted edge existence scores. The AUROCs are very high when we compare the positive transactions to all negatives, and while it drops as $t$ increases, it remains above 0.93 even seven timesteps after the end of the training data. As expected, the AUROCs are lower when we only include negatives that appeared in the training data, since those are harder to discriminate from the true positives. However, the AUROCs still remain decently high for the train negatives, especially for the first few timesteps. This ability to accurately generate transactions under shocks, especially for multiple future timesteps, could help decision-makers forecast which firms or products will be affected next, enable early detection of incoming disruptions, and support sophisticated simulations of interventions.
6 Discussion